OpenStudy (anonymous):

In the equation 8(Ax-7)-5(2x-1)=14x+B, how does B=-51?

OpenStudy (anonymous):

open brackets 8Ax-7*8 - 10x +5 =14x + B (8A-10)X -51 = 14X +B Compare B= -51

OpenStudy (cggurumanjunath):